find two numbers whose sum is 55 and whose difference is 15 let the numbers be x and y thank you
Answer by checkley75(3666)

x+y=55
x-y=15 or x=15+y
(15+y)+y=55
15+2y=55
2y=55-15
2y=40
y=40/2
y=20 ans.
x+20=55
x=55-20
x=35 ans.